# PolicyPad: Collaborative Prototyping of LLM Policies

### v1

@@ Appendix C Themes from Formative Study Qualitative Coding
Table 2. Our 6 themes that emerged from an analysis of transcripts from our observational study.:
Appendix D Findings and System Iterations from Co-design Sessions
D.1. Version 1
@@
Figure 11. Version 1 of PolicyPad: a simple collaborative policy editor with a policy-informed model in the sidebar.:
@@
The goal for our initial version of the system (Fig 11) was simplicity: we wanted to validate the core premise of our workflow before adding complexity. We built a basic collaborative document editor [^10] We showed participants alternative editors besides documents, such as a node-based interfaces (Arawjo et al., 2024; Pu et al., 2025), as a design exploration, but they found them too unfamiliar and unnecessarily complex. with an LLM chatbot in a sidebar that used the contents of the document as its policy. Contents of the document are shared across all users whereas the sidebar is for personal experimentation. Scenarios and heuristics were provided to participants in a separate Google Doc. Participants appreciated the collaborative nature of the document and easy access to the policy-informed model, which allowed them to quickly iterate on the policy. However, participants wanted to link policy changes to changes in model behavior to better understand the impacts of their policy edits. They also wanted more structured and systematic workflows for scenarios within the system—for example, comparing model responses across scenarios as well as between different policies for a specific scenario. Finally, the policy editor was a bit too simple, and participants wanted richer editing and formatting support.
D.2. Version 2
@@
Figure 12. Version 2 of PolicyPad: a block-based editor with policy versioning and more support for structured interaction with scenarios in the sidebar. The top screen shows the sidebar in a collapsed state. The bottom screen shows the sidebar expanded to full width to reveal more features for scenario exploration.:
@@
The second version of the system (Fig. 12) featured a block editor (similar to Notion) with expressive editing and formatting functionality. We added a persistent right side panel with a “scenario gallery” that allows users to explore scenarios (a user query followed by an AI response) and stress-test the policy by extending the conversation. We also introduced policy versioning, as well AI-generated notes summarizing 1) the nature of the policy update, and 2) changes to the response to a particular scenario due to the policy update. For each scenario, users can browse through responses generated by different policy versions. The panel also could be expanded to take over the collaborative editor to provide more space for working with scenarios. Overall, participants thought this version was a significant improvement over the previous one. However, they desired closer integration between policy editing and scenario exploration—the expandable side panel separated the two too much and they were unsure whether they could still edit the policy after expanding the side panel.
D.3. Version 3
@@
Figure 13. Version 3 of PolicyPad: we used the same block-based editor as Version 2 but with more closely integrated scenarios into the collaborative policy editor via interactive pill-shaped widgets.:
@@
In the third version (Fig 13), we removed the persistent right side panel and represented scenarios as interactive widgets within the policy editor itself to tighten the relationship between policy editing and scenario exploration. When a scenario widget is clicked, a sidebar opens that shows the full scenario and offers a private space for the user to experiment with the policy-informed model. Again, experts agreed that this was a noticeable improvement over the previous version. However, because the sidebar is private, they suggested adding features that would allow users to flag or share specific scenarios or responses with the broader group for discussion. They also viewed notes summarizing policy and response changes as potentially unnecessary to reduce clutter in the sidebar. We incorporated this feedback into the final design of our system.
Appendix E Starter Heuristics and Policy Objectives Section
@@
Appendix F Post-Study Policy Rating Questions
@@
Appendix G Evaluation Study Participants
@@
Table 3. Details of participants (gender, age range, education status, years of practical experience, and generative AI use) in our evaluation study. The “GenAI use” column refers to participants’ experience using generative AI tools, whether it be personally or professionally, as determine by their frequency of use. The response “Occasional” corresponds to the following description: “I’ve tried it here and there but don’t use it regularly.” All participants specializing in mental health were based in the U.S.. All participants specializing law except two (who were based in Europe and Asia, respectively) were based in the U.S..:

### v3

@@ Appendix A Themes from Formative Study Qualitative Coding
Table 1. Our 6 themes that emerged from an analysis of transcripts from our observational study.
A 2-column table with headings Theme and Description. There are 6 rows with the 6 high-level themes alongside their descriptions.
Appendix B Findings and System Iterations from Co-design Sessions
B.1. Version 1
@@
Figure 12. Version 1 of PolicyPad: a simple collaborative policy editor with a policy-informed model in the sidebar.
A UI showing Version 1 of PolicyPad: a simple collaborative policy editor with a policy-informed model in the sidebar
@@
The goal for our initial version of the system (Fig 12) was simplicity: we wanted to validate the core premise of our workflow before adding complexity. We built a basic collaborative document editor [^25] We showed participants alternative editors besides documents, such as a node-based interfaces (Arawjo et al., 2024; Pu et al., 2025), as a design exploration, but they found them too unfamiliar and unnecessarily complex. with an LLM chatbot in a sidebar that used the contents of the document as its policy. At first, the document was blank—experts wrote the policy from scratch via the activity described in the “Co-design workshop 1” section of our Supplementary Materials.
The policy (i.e, contents of the document) are shared across all users whereas the sidebar is for personal experimentation. Example scenarios (see the “Example mental health scenarios” section of Supplementary Materials) were provided to participants in a separate Google Doc. Participants appreciated the collaborative nature of the document editor and easy access to the policy-informed model, [^26] Recall from the paper that the policy-informed model is an LLM that has the policy incorporated into its system instructions such that its behavior is informed by the policy. which allowed them to quickly iterate on the policy. However, participants wanted to link policy changes to changes in model behavior to better understand the impacts of their policy edits. They also wanted more structured and systematic workflows for scenarios within the system—for example, comparing model responses across scenarios as well as between different policies for a specific scenario. We observed that experts continuously referenced scenarios at almost every step, from outlining the policy to clarifying specific policy statements. This validated the importance designing for structured interaction with scenarios and smooth integration of scenarios into policy editing workflows. Finally, the policy editor was a bit too simple, and participants wanted richer editing and formatting support.
B.2. Version 2
@@
Figure 13. Version 2 of PolicyPad: a block-based editor with policy versioning and more support for structured interaction with scenarios in the sidebar. The top screen shows the sidebar in a collapsed state. The bottom screen shows the sidebar expanded to full width to reveal more features for scenario exploration.
A UI showing Version 2 of PolicyPad: a block-based editor with policy versioning and more support for structured interaction with scenarios in the sidebar. The top screen shows the sidebar in a collapsed state. The bottom screen shows the sidebar expanded to full width to reveal more features for scenario exploration
@@
The second version of the system (Fig. 13) featured a block editor (similar to Notion) with expressive editing and formatting functionality. We added a persistent right side panel with a “scenario gallery” that allows users to explore scenarios (a user query followed by an AI response) and stress-test the policy by extending the conversation. We also introduced policy versioning, as well AI-generated notes summarizing 1) the nature of the policy update, and 2) changes to the response to a particular scenario due to the policy update. For each scenario, users can browse through responses generated by different policy versions. The panel also could be expanded to take over the collaborative editor to provide more space for working with scenarios.
Overall, participants thought this version was a significant improvement over the previous one. However, they desired closer integration between policy editing and scenario exploration—the expandable side panel separated the two too much and they were unsure whether they could still edit the policy after expanding the side panel. E9, for example, shared: “[in Session 1] we were able to write a rule and see how that changed the response. That was one of the key parts. Is that something you can do in this one?” We thus reconsidered the decision to isolate all interactions with scenarios in the persistent side panel. The facilitator’s guide for this co-design session can be found in the “Co-design workshop 2” section of our Supplementary Materials.
B.3. Version 3
@@
Figure 14. Version 3 of PolicyPad: we used the same block-based editor as Version 2 but with more closely integrated scenarios into the collaborative policy editor via interactive pill-shaped widgets.
A UI showing Version 3 of PolicyPad: the same block-based editor was used compared to Version 2 but with more closely integrated scenarios into the collaborative policy editor via interactive pill-shaped widgets.
@@
In the third version (Fig 14), we removed the persistent right side panel and represented scenarios as interactive widgets within the policy editor itself to tighten the relationship between policy editing and scenario exploration. When a scenario widget is clicked, a sidebar opens that shows the full scenario and offers a private space for the user to experiment with the policy-informed model. Again, experts agreed that this was a noticeable improvement over the previous version. However, because the sidebar is private, many (including E7) suggested adding features that would allow users to flag or share specific scenarios or responses with the broader group for discussion. We also provided an explicit save button that controlled when a new model response version was generated, which experts appreciated: “I’d rather have to save the policy to dictate when I get to do the comparison.” [E4]. They also viewed notes summarizing policy and response changes as potentially unnecessary to reduce clutter in the sidebar. We incorporated this feedback into the final design of our system. The facilitator’s guide for this co-design session can be found in the “Co-design workshop 3” section of our Supplementary Materials.
Appendix C Connections Between UX and LLM Policy Prototyping
See Table 2.
Observation Relevant UX Method UX Definition & Usage Usage in LLM Policy Prototyping PolicyPad Features
Incomplete feedback loops (Section 3.2.1) Rapid Prototyping (e.g., (Lim et al., 2008; Houde and Hill, 1997; Buchenau and Suri, 2000; Camburn et al., 2017; Dow et al., 2010; Quicksey and Meierling, 2022; Kontschieder, 2018)) Tight feedback loops of ideating, implementing, and evaluating design ideas. Allows designers to identify usability issues early, explore alternatives, and align teams to shared visions Tight feedback loops of ideating, drafting, and testing policy statements for quick identification of policy “usability” issues (e.g., unclear statements), characteristics of responsible model behavior, and translations of that behavior into policy. Response (re)generation with policy-informed model, policy suggestion upon editing response.
Operating at both high and low levels (Section 3.2.2) Low-fidelity prototyping (e.g., (Rudd et al., 1996; Virzi et al., 1996; Walker et al., 2002)) Artifact loosely resembling the final product in terms of look & feel and/or implementation. Cheap to create and iterate upon, ideal for collecting early requirements and feedback. Artifact providing high-level documentation of responsible model behavior to quickly gather and integrate perspectives/feedback. Requires focus on high-level details. We focus on this type of policy prototype in our work. Heuristics editor and checker, freeform document editor.
Operating at both high and low levels (Section 3.2.2) High-fidelity prototyping (e.g., (Rudd et al., 1996; Virzi et al., 1996; Walker et al., 2002)) Artifact closely resembling the final product in terms of look & feel and/or implementation. They are useful for collecting detailed feedback but may be expensive to create. Artifact providing detailed documentation of responsible model behavior to guide alignment efforts, often with polished wording, illustrative examples, legal sign-off, and more. Requires focus on high- and low-level details. Example: OpenAI Model Spec (OpenAI, 2025b). N/A—focus of PolicyPad is low-fidelity prototyping.
Scenarios grounded discussions (Section 3.2.3) Storyboarding/ scenario-building (e.g., (Figma, 2025; Truong et al., 2006; Hooper and Hsia, 1982; Bodker, 1999; Andersen and Jæger, 1999)) Concrete representations of users, contexts, and tasks to ground abstract design ideas. Panels add context and illustrate user stories. Promotes reflection and communication among stakeholders. Sample user-AI conversations to ground policy discussions and creation. Conversational turns add context and illustrate sample user & model behaviors. Promotes reflection and communication among stakeholders. Interactive scenarios, adding/extending scenarios, spotlight scenarios.
Experts valued synchronous collaboration (Section 3.2.4) Design workshops e.g., (Rosner et al., 2016; Elsden et al., 2020; Rosenbaum et al., 2002; Kimbell and Bailey, 2017; Hagan, 2021) Common collaborative method for gathering user requirements, studying empirical phenomena, and evaluating interactive systems. Can serve as a field site, research instrument, or a research account. Small-group sessions that serve as a field site for collective ideation and reflection of responsible model behavior in domain-specific use cases. Collaborative multi-player editor, spotlight scenarios, response flagging.
Table 2. Mapping of UX methods relevant to insights from our observational study (Section 3) to their usage in LLM policy prototyping, to features in PolicyPad supporting that usage.
A 5-column table with 5 rows, mapping UX concepts of rapid prototyping, low-fidelity prototyping, high-fidelity prototyping, storyboarding, and design workshops to their usage in LLM policy prototyping.
Appendix D Starter Heuristics and Policy Objectives Section
@@
Appendix E Post-Study Policy Rating Questions
@@
Appendix F Evaluation Study Participants
@@
Table 3. Details of participants (gender, age range, education status, years of practical experience, and generative AI use) in our evaluation study. The “GenAI use” column refers to participants’ experience using generative AI tools, whether it be personally or professionally, as determine by their frequency of use. The response “Occasional” corresponds to the following description: “I’ve tried it here and there but don’t use it regularly.” All participants specializing in mental health were based in the US. All participants specializing law except two (who were based in Europe and Asia, respectively) were based in the US.
A table summarizing the background information of participants (P1 to P22) grouped into eight groups in our evaluation study.
Appendix G Areas of Disagreement Across Expert Groups